OpenStudy (anonymous):

Please help I'm really stuck on these two problems Simplify 3 ^4√3 - 8 ^3√3 + 5 ^ 4√3 + 4 ^3√3 a.2√2 b.-5√40 c.3√14-2√6 d.3√28-4√3 Which of the following represents ^9√x^7 in exponential form? my options are a.x^9/7 b.x^7/9 c.9x^7 d.7x^9 Thank you so much for your help in advance.

OpenStudy (anonymous):

i don't understand what "3 ^4√3" means

OpenStudy (anonymous):

it means that the 4 is on top of that triangle part of the radical sign

OpenStudy (anonymous):

\[3\sqrt[4]{3}-8\sqrt[3]{3}+5\sqrt[4]{3}+4\sqrt[3]{3}\]

OpenStudy (anonymous):

combine like terms, like in \[3x-8y+5x+4y\]

OpenStudy (anonymous):

you get \[8\sqrt[4]{3}-4\sqrt[3]{3}\]

OpenStudy (anonymous):

Thank you so much!